# Learning Commonality, Divergence and Variety for Unsupervised Visible-Infrared Person Re-identification

### v2

@@ Abstract
Unsupervised visible-infrared person re-identification (USVI-ReID) aims to match specified people in infrared images to visible images without annotation, and vice versa. USVI-ReID is a challenging yet under-explored task. Most existing methods address the USVI-ReID problem using cluster-based contrastive learning, which simply employs the cluster center as a representation of a person. However, the cluster center primarily focuses on commonality, overlooking divergencies. To address the problem, we propose a Progressive Contrastive Learning with Hard and Dynamic Prototypes method for USVI-ReID. In brief, we first generate the hard prototype by selecting the sample with the maximum distance from the cluster center. We theoretically show that the hard prototype is used in the contrastive loss to emphasize divergence. Additionally, instead of rigidly aligning query images to a specific prototype, we generate the dynamic prototype by randomly picking samples within a cluster. The dynamic prototype is used to encourage the variety of features. Finally, we introduce a progressive learning strategy to gradually shift the model’s attention towards divergence and variety, avoiding cluster deterioration. Extensive experiments conducted on the publicly available SYSU-MM01 and RegDB datasets validate the effectiveness of the proposed method. Our method outperforms the existing state-of-the-art method with an average mAP improvement of 3.9%. The source codes will be released.
@@ 1 Introdction
Visible-infrared person re-identification (VI-ReID) aims at matching the same person captured in one modality with their counterparts in another modality [1, 2, 3, 4]. It has recently gained attention in computer vision applications like video surveillance, security, and person analysis. With the development of deep learning, VI-ReID has achieved remarkable advancements [5, 6, 7, 8, 9]. However, the development of existing VI-ReID methods is still limited due to the requirement for expensive-annotated training data [10]. To mitigate the problem of annotating large-scale cross-modality data, some semi-supervised VI-ReID methods [11, 12, 13] are proposed to learn modality-invariant and identity-related discriminative representations by utilizing both labeled and unlabeled data. For this purpose, OTLA [11] proposed an optimal transport label assignment mechanism to assign pseudo-labels for unlabeled infrared images while ignoring how to calibrate noise pseudo-labels. DPIS [12] integrates two pseudo-labels generated by distinct models into a hybrid pseudo-label for unlabeled infrared data, but it makes the training process more complex. Although these methods have gained promising performances, they still rely on a certain number of manual-labeled data.
@@
Several USVI-ReID methods [14, 15, 16, 17] have proposed to tackle the issues of expensive visible-infrared annotation through contrastive learning. These methods create two modality-specific memories, one for visible features and the other for infrared features. During training, these methods consider the memory center as a prototype and minimize the contrastive loss across the features of query images and prototype. Then, these methods aggregate the corresponding prototypes based on similarity. However, the centroid prototype only stores the commonality of each person, neglecting the divergence, which causes the pseudo-labels generated by the cluster to be unreliable. Just like a normal distribution, to better reflect the data distribution of a dataset, we need not only the mean but also the variance. To this end, we pay attention to the cluster divergence, constraining the distance between the furthest points and the centroid to ensure cluster compactness.
@@
In this paper, we argue that an important aspect of contrastive learning for USVI-ReID, i.e. the design of the prototype, has so far been neglected, and propose progressive contrastive learning with hard and dynamic prototype (PCLHD) method for the USVI-ReID. Firstly, we design a Hard Prototype Contrastive Learning (HPCL) to mine divergent yet meaningful information. In contrast to traditional contrastive learning methods, we choose the hard samples to serve as the hard prototype. In other words, the hard prototype is the one that is farthest from the memory center. The hard prototype encompasses distinctive information. Furthermore, we introduce the concept of Dynamic Prototype Contrastive Learning (DPCL), we randomly select samples from each cluster to serve as the dynamic prototype. DPCL effectively accounts for the intrinsic variety within clusters, enhancing the model’s adaptability to varying data distributions. Early clustering results are unreliable, and utilizing hard samples at this stage may lead to cluster degradation. Therefore, we introduce progressive contrastive learning to gradually focus on hard samples.
@@
  • Experiments on SYSU-MM01 and RegDB datasets demonstrate the superiority of our method compared with existing USVI-ReID methods, and PCLHD generates higher-quality pseudo-labels than other methods.
@@ 2.1 Supervised Visible-Infrared Person ReID
Visible-Infrared Person ReID (VI-ReID) has drawn much attention in recent years [1, 18, 19, 20, 21, 22]. Many VI-ReID works focused on mitigating huge semantic gaps across modalities have made advanced progress, which can be classified into two primary classes based on their different aligning ways: image-level alignment and feature-level alignment. The image-level alignment methods focus on reducing cross-modality gaps by modality translation. Some GAN-based methods [23, 24] are proposed to perform style transformation for aligning cross-modality images. However, the generated images unavoidably contain noise. Therefore, X-modality [25] and its promotion [26, 27] align cross-modality images by introducing a middle modality. Mainstream feature-level alignment methods [28, 29, 30, 31, 32] focus on minimizing cross-modality gaps by finding a modality-shared feature space. However, the advanced performances of the above methods build on large-scale human-labeled cross-modality data, which are quite time-consuming and expensive, thus hindering the fast application of these methods in real-scenes.
@@ 2.2 Unsupervised Single-Modality Person ReID
The existing unsupervised single-modality person ReID methods can be roughly divided into two classes: Unsupervised domain adaption methods (UDA), which try to leverage the knowledge transferred from labeled source domain to improve performance [33, 34, 35, 36, 37], and fully unsupervised methods (USL), which directly train a USL-ReID model on the unlabeled target domain [15, 38]. Compared with the UDA methods, the USL methods are more challenging. Recently, cluster-contrast learning [39] has achieved impressive performance by performing contrastive learning at the cluster level. However, cluster-contrast with a uni-proxy can be biased and confusing, which fails to accurately describe the information of a cluster. To this end, methods [40, 41] proposed maintaining multi-proxies for a cluster to adaptively capture different information within the cluster. The above methods are mainly proposed to solve the single-modality ReID task, but they are limited to solving the USL-VI-ReID task due to large cross-modality gaps.
@@ 2.3 Unsupervised Visible-Infrared Person ReID
Unsupervised visible-infrared person ReID has attracted much attention due to the advantage of not relying on any data annotation. Some UDA methods [42, 11] use a well-annotated labeled source domain for pre-training to solve the USVI-ReID task. Some fully unsupervised methods [14, 17, 16, 43] adopt contrastive learning to boost performance inspired by [39], which mainly follow a two-step loop paradigm: generating pseudo-labels using the DBSCAN algorithm [44] to create memory banks with clustering centers and establishing cross-modality correspondences based on these memory banks. However, pseudo-labels are often inaccurate and rigid, CCLNet [45] leverages the text information from CLIP to afford greater semantic monitoring insights to compensate for the rigidity of pseudo-labels. Moreover, reliable cross-modality correspondences are vital to USVI-ReID, thus PGM [17] proposes a progressive graph matching framework to establish more reliable cross-modality correspondences. However, cluster centers mainly present common information while lacking distinctive information, which results in ambiguous cross-modality correspondences when meeting hard samples[46, 47].
@@ 3.1 Problem Formulation and Overview
We propose a Progressive Contrastive Learning with Hard and Dynamic Prototype (PCLHD) method for USVI-ReID, which mainly contains an online encoder, momentum encoder, and progressive contrastive learning strategy with a centroid prototype, a hard prototype, and a dynamic prototype, as shown in Fig. 1. The online encoder is a standard network, updated through back-propagation. The momentum encoder mirrors the structure of the online encoder, updated through the weights of the online encoder. The clustering is used to generate pseudo labels for creating cluster-aware memory, and we employ DBSCAN [44] for clustering. PCLHD primarily focuses on representation learning, and we use PGM [17] to aggregate cross-modality memory.
@@ 3.2 Centroid Prototype Contrastive Learning
Following the USVI-ReID methods [14, 16, 48], we use centroid prototype contrastive learning to optimize the online encoder in the first state, which includes memory initialization and optimization.
@@ 3.3 Hard Prototype Contrastive Learning
To ensure that the prototype effectively captures divergence within a identity, we devise a novel hard prototype for contrastive learning, which is referred to as Hard Prototype Contrastive Learning. HPCL is designed to provide a comprehensive understanding of personal characteristics, which benefits its handling of hard samples. We use the online encoder $\phi_{0}$ to extract feature representations, and select $k$ samples that are farthest from the memory center as the hard prototype:
@@
Theorem The information entropy of hard sample prototypes is greater than the information entropy of centroid prototypes, thereby preserving greater divergence within the hard memory.
Proof Given a set of features $\{\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{N_{c}}\}$ for class $c$. The entropy $H(\mathbf{cm}_{c})$ can be approximated by the entropy of the distribution of the sample means. Considering that $\mathbf{cm}_{c}$ is a convex combination of the sample features $\mathbf{z}_{i}$, we have:
@@ 3.4 Dynamic Prototype Contrastive Learning
Inspired by MoCo [49], we design dynamic prototype contrastive learning in order to preserve the intrinsic variety in sample features. DPCL comprises an online encoder $\phi_{0}$ and a momentum encoder $\phi_{m}$. The momentum encoder mirrors the structure of the online encoder, which is updated by the accumulated weights of the online encoder:
@@ 4.1 Experiment Setting
Implementation Details.  We adopt the feature extractor in AGW [51], which is initialized with ImageNet-pretrained weights to extract 2048-dimensional features. During the training stage, the input images are resized to 288$\times$144. We follow augmentations in  [1] for data augmentation. In one batch, we randomly sample 16 person IDs, and each person ID samples 16 instances. We set $M$ to be 16 for computational convenience. The number of epochs is 100, in which the first 50 epochs are trained by contrastive loss with the centroid prototype. For the last 50 epochs, we train the model by contrastive loss with both the hard and dynamic prototypes. $E_{CPCL}$ is 50. At the beginning of each epoch, we utilize the DBSCAN [44] algorithm to generate pseudo labels. During the inference stage, we use the momentum encoder $\phi_{m}$ to extract features and take the features of the global average pooling layer to calculate cosine similarity for retrieval. The momentum value $\alpha$ and $\beta$ is set to 0.1 and 0.999, respectively. The temperature hyper-parameter $\tau$ is set to 0.05 and the weighting hyper-parameter $\lambda$ in Eq.(23) is 0.5.
@@ 4.2 Results and Analysis
To comprehensively evaluate our method, we compare our method with 18 supervised VI-ReID methods, 3 semi-supervised VI-ReID methods, and 9 unsupervised VI-ReID methods. The comparison results on the SYSU-MM01 and RegDB are reported in Tab. 1.
@@
Comparison with SSVI-ReID Methods. There are three SSVI-ReID methods proposed to alleviate the problem of labeling cost by using a part of data annotations. Remarkably, our method’s performance without any annotations surpasses that of all methods relying on partial annotation.
Comparison with SVI-ReID Methods. Surprisingly, our method without annotation outperforms several SVI-ReID methods, e.g., DDAG [30], AGW [51], NFS [52], LbA [53]. This shows the immense competitiveness of PCLHD compared to SVI-ReID methods that rely on complete data labeling. The superior performance of PCLHD mainly benefits from the hard prototype and dynamic prototype contrastive learning. Additionally, it’s worth acknowledging that there is still a significant disparity between PCLHD and the state-of-the-art fully-supervised results.
@@ 4.4 Further Analysis
The ARI metric. Following MMM [48], we utilize the Adjusted Rand Index (ARI) metric for clustering evaluation. The larger the ARI value, the higher the clustering quality. In Fig. 2 (c), “RGB” and “IR” denote the ARI values of visible and infrared clusterings, which can measure the quality of visible and infrared pseudo-labels. “ALL” means the ARI values of overall clusterings, which can evaluate the reliability of cross-modality correspondences. PCLHD surpasses other methods significantly on all of the mentioned ARI values, which demonstrates our method can effectively mine divergence and variety to improve clustering quality.

### v3

@@ Abstract
Unsupervised visible-infrared person re-identification (USVI-ReID) aims to match specified people in infrared images to visible images without annotations, and vice versa. USVI-ReID is a challenging yet under-explored task. Most existing methods address the USVI-ReID using cluster-based contrastive learning, which simply employs the cluster center as a representation of a person. However, the cluster center primarily focuses on commonality, overlooking divergence and variety. To address the problem, we propose a Progressive Contrastive Learning with Hard and Dynamic Prototypes method for USVI-ReID. In brief, we generate the hard prototype by selecting the sample with the maximum distance from the cluster center. We theoretically show that the hard prototype is used in the contrastive loss to emphasize divergence. Additionally, instead of rigidly aligning query images to a specific prototype, we generate the dynamic prototype by randomly picking samples within a cluster. The dynamic prototype is used to encourage the variety. Finally, we introduce a progressive learning strategy to gradually shift the model’s attention towards divergence and variety, avoiding cluster deterioration. Extensive experiments conducted on the publicly available SYSU-MM01 and RegDB datasets validate the effectiveness of the proposed method.
@@ 1 Introdction
Visible-infrared person re-identification (VI-ReID) aims at matching the same person captured in one modality with their counterparts in another modality [1, 2, 3, 4]. It has recently gained attention in computer vision applications like video surveillance [5] and image retrieval [6, 7]. With the development of deep learning [8, 9, 10, 11], VI-ReID has achieved remarkable advancements [12, 13, 14]. However, the development of existing VI-ReID methods is still limited due to the requirement for expensive-annotated training data [15, 16]. To mitigate the problem of annotating large-scale cross-modality data, some semi-supervised VI-ReID methods [17, 18, 19] are proposed to learn modality-invariant and identity-related discriminative representations by utilizing both labeled and unlabeled data. For this purpose, OTLA [17] proposed an optimal transport label assignment mechanism to assign pseudo-labels for unlabeled infrared images while ignoring how to calibrate noise pseudo-labels. DPIS [18] integrates two pseudo-labels generated by distinct models into a hybrid pseudo-label for unlabeled infrared data, but it makes the training process more complex. Although these methods have gained promising performances, they still rely on a certain number of manual-labeled data.
@@
Several USVI-ReID methods [20, 21, 22, 23] have proposed to tackle the issues of expensive visible-infrared annotation through contrastive learning. These methods create two modality-specific memories, one for visible features and the other for infrared features. During training, these methods consider the memory center as a prototype and minimize the contrastive loss across the features of query images and prototype. Then, these methods aggregate the corresponding prototypes based on similarity. However, the centroid prototype only stores the commonality of each person, neglecting the divergence [24, 25, 26], which causes the pseudo-labels generated by the cluster to be unreliable. Just like a normal distribution, to better reflect the data distribution of a dataset, we need not only the mean but also the variance.
@@
In this paper, we argue that an important aspect of contrastive learning for USVI-ReID, i.e. the design of the prototype, has so far been neglected, and propose progressive contrastive learning with hard and dynamic prototype (PCLHD) method for the USVI-ReID. Firstly, we design a Hard Prototype Contrastive Learning (HPCL) to mine divergent yet meaningful information. In contrast to traditional contrastive learning methods, we choose the hard samples to serve as the hard prototype. In other words, the hard prototype is the one that is farthest from the memory center. The hard prototype encompasses distinctive information. Furthermore, we introduce the concept of Dynamic Prototype Contrastive Learning (DPCL), we randomly select samples from each cluster to serve as the dynamic prototype. DPCL effectively accounts for the intrinsic variety within clusters, enhancing the model’s adaptability to varying data distributions. Early clustering results are unreliable, and utilizing hard and dynamic prototype at this stage may lead to cluster degradation. Therefore, we introduce progressive contrastive learning to gradually focus on divergence and variety.
@@
  • Experiments on SYSU-MM01 and RegDB datasets demonstrate the superiority of our method compared to existing USVI-ReID methods, and PCLHD generates higher-quality pseudo-labels than other methods.
@@ 2.1 Supervised Visible-Infrared Person ReID
Visible-infrared person re-identification (VI-ReID) has drawn much attention in recent years [27, 28, 29, 30, 31, 32]. Many VI-ReID works focused on mitigating huge semantic gaps across modalities have made advanced progress, which can be classified into two primary classes based on their different aligning ways: image-level alignment and feature-level alignment. The image-level alignment methods focus on reducing cross-modality gaps by modality translation. Some GAN-based methods [33, 34] are proposed to perform style transformation for aligning cross-modality images. However, the generated images unavoidably contain noise. Therefore, X-modality [35] and its promotions [36, 37] align cross-modality images by introducing a middle modality. Mainstream feature-level alignment methods [38, 39, 40] focus on minimizing cross-modality gaps by finding a modality-shared feature space. However, the advanced performances of the above methods build on large-scale human-labeled cross-modality data, which are quite time-consuming and expensive, thus hindering the fast application of these methods in real-scenes.
@@ 2.2 Unsupervised Single-Modality Person ReID
The existing unsupervised single-modality person ReID methods can be roughly divided into two classes: Unsupervised domain adaption (UDA) methods, which try to leverage the knowledge transferred from labeled source domain to improve performance [41, 42, 43, 44], and fully unsupervised methods (USL), which directly train a USL-ReID model on the unlabeled target domain [21, 45]. Compared with the UDA methods, the USL methods are more challenging. Recently, cluster-contrast learning [46] has achieved impressive performance by performing contrastive learning at the cluster level. However, cluster-contrast with a uni-proxy can be biased and confusing, which fails to accurately describe the information of a cluster. To this end, the methods [47, 48] proposed maintaining multi-proxies for a cluster to adaptively capture different information within the cluster. The above methods are mainly proposed to solve the single-modality ReID task, but they are limited to solving the USL-VI-ReID task due to large cross-modality gaps.
@@ 2.3 Unsupervised Visible-Infrared Person ReID
Unsupervised visible-infrared person ReID (USVI-ReID) has attracted much attention due to the advantage of not relying on any data annotation. Some UDA methods [49, 17] use a well-annotated labeled source domain for pre-training to solve the USVI-ReID task. Some fully unsupervised methods [23, 22] adopt contrastive learning to boost performance, which mainly follow a two-step loop paradigm: generating pseudo-labels using the DBSCAN algorithm [50] to create memory banks with clustering centers and establishing cross-modality correspondences based on these memory banks. However, pseudo-labels are often inaccurate and rigid, CCLNet [51] leverages the text information from CLIP to afford greater semantic monitoring insights to compensate for the rigidity of pseudo-labels. Moreover, reliable cross-modality correspondences are vital to USVI-ReID, thus PGM [23] proposes a progressive graph matching framework to establish more reliable cross-modality correspondences. However, cluster centers mainly present common information while lacking distinctive information, which results in ambiguous cross-modality correspondences when meeting hard samples [52, 53].
@@ 3.1 Problem Formulation and Overview
We propose a Progressive Contrastive Learning with Hard and Dynamic Prototype (PCLHD) method for USVI-ReID, which mainly contains online encoder, momentum encoder, and progressive contrastive learning strategy with centroid prototype, hard prototype, and dynamic prototype, as shown in Fig. 1. The online encoder is a standard network, updated through back-propagation. The momentum encoder mirrors the structure of the online encoder, updated through the weights of the online encoder. The clustering is used to generate pseudo labels for creating cluster-aware memory, and we employ DBSCAN for clustering. PCLHD primarily focuses on representation learning, and we use PGM [23] to aggregate cross-modality memory.
@@ 3.2 Centroid Prototype Contrastive Learning
Following the USVI-ReID methods [22, 54], we use centroid prototype contrastive learning to optimize the online encoder in the first state, which includes memory initialization and optimization.
@@ 3.3 Hard Prototype Contrastive Learning
To ensure that the prototype effectively captures divergence within a identity, we devise a novel hard prototype for contrastive learning, which is referred to as Hard Prototype Contrastive Learning. HPCL is designed to provide a comprehensive understanding of personal characteristics, which benefits its handling of hard samples [47]. We use the online encoder $\phi_{0}$ to extract feature representations, and select $k$ samples that are farthest from the memory center as the hard prototype:
@@
Theorem 1. The information entropy of hard sample prototypes is greater than the information entropy of centroid prototypes, thereby preserving greater divergence within the hard memory.
Given a set of features $\{\mathbf{z}_{1},\mathbf{z}_{2},\ldots,\mathbf{z}_{N_{c}}\}$ for class $c$. The entropy $H(\mathbf{cm}_{c})$ can be approximated by the entropy of the distribution of the sample means. Considering that $\mathbf{cm}_{c}$ is a convex combination of the sample features $\mathbf{z}_{i}$, we have:
@@ 3.4 Dynamic Prototype Contrastive Learning
Inspired by MoCo [55] and DPM [56], we design dynamic prototype contrastive learning in order to preserve the intrinsic variety in sample features. DPCL comprises an online encoder $\phi_{0}$ and a momentum encoder $\phi_{m}$. The momentum encoder mirrors the structure of the online encoder, which is updated by the accumulated weights of the online encoder:
@@ 4.1 Experiment Setting
Implementation Details.  We adopt the feature extractor in AGW [58], which is initialized with ImageNet-pretrained weights to extract 2048-dimensional features. During the training stage, the input images are resized to 288$\times$144. We follow augmentations in CAJ  [1] for data augmentation. In one batch, we randomly sample 16 pseudo identities, and each pseudo identity samples 16 instances. We set $M$ to be 16 for computational convenience. The number of epochs is 100, in which the first 50 epochs are trained by contrastive loss with the centroid prototype. For the last 50 epochs, we train the model by contrastive loss with both the hard and dynamic prototypes. $E_{CPCL}$ is 50. At the beginning of each epoch, we utilize the DBSCAN [50] algorithm to generate pseudo labels. During the inference stage, we use the momentum encoder $\phi_{m}$ to extract features and take the features of the global average pooling layer to calculate cosine similarity for retrieval. The momentum value $\alpha$ and $\beta$ is set to 0.1 and 0.999, respectively. The temperature hyper-parameter $\tau$ is set to 0.05 and the weighting hyper-parameter $\lambda$ in Eq.(23) is 0.5.
@@ 4.2 Results and Analysis
To comprehensively evaluate our method, we compare our method with 19 supervised VI-ReID methods, 3 semi-supervised VI-ReID methods, and 9 unsupervised VI-ReID methods. The comparison results on the SYSU-MM01 and RegDB are reported in Tab. 1.
@@
Comparison with SSVI-ReID Methods. There are three SSVI-ReID methods proposed to alleviate the problem of labeling cost by using a part of annotations. Remarkably, our method achieves superior performance without any annotations, outperforming all existing SSVI-ReID methods that utilize partial annotations. Moreover, the results suggest that our method can significantly reduce the dependency on manual annotations.
Comparison with SVI-ReID Methods. Surprisingly, our method without annotation outperforms several SVI-ReID methods, e.g., DDAG [39], AGW [58], NFS [59], LbA [60]. This shows the immense competitiveness of PCLHD compared to SVI-ReID methods that rely on complete data annotations. The superior performance of PCLHD mainly benefits from the hard prototype and dynamic prototype contrastive learning. Additionally, we have to acknowledge that a significant disparity still exists between PCLHD and the state-of-the-art fully-supervised results.
@@ 4.4 Further Analysis
The ARI metric. Following MMM [54], we utilize the Adjusted Rand Index (ARI) metric for clustering evaluation. The larger the ARI value, the higher the clustering quality. In Fig. 2 (c), “RGB” and “IR” denote the ARI values of visible and infrared clusterings, which can measure the quality of visible and infrared pseudo-labels. “ALL” means the ARI values of overall clusterings, which can evaluate the reliability of cross-modality correspondences. PCLHD surpasses other methods significantly on all of the mentioned ARI values, which demonstrates PCLHD can effectively mine divergence and variety to improve clustering quality.